# Soft Prompting for Unlearning in Large Language Models

### v1

@@ Abstract
The widespread popularity of Large Language Models (LLMs), partly due to their unique ability to perform in-context learning, has also brought to light the importance of ethical and safety considerations when deploying these pre-trained models. In this work, we focus on investigating machine unlearning for LLMs motivated by data protection regulations. In contrast to the growing literature on fine-tuning methods to achieve unlearning, we focus on a comparatively lightweight alternative called soft prompting to realize the unlearning of a subset of training data. With losses designed to enforce forgetting as well as utility preservation, our framework Soft Prompting for Unlearning (SPUL) learns prompt tokens that can be appended to an arbitrary query to induce unlearning of specific examples at inference time without updating LLM parameters. We conduct a rigorous evaluation of the proposed method and our results indicate that SPUL can significantly improve the trade-off between utility and forgetting in the context of text classification with LLMs. We further validate our method using multiple LLMs to highlight the scalability of our framework and provide detailed insights into the choice of hyperparameters and the influence of the size of unlearning data. Our implementation is available at https://github.com/karuna-bhaila/llm_unlearning.
@@ 1 Introduction
In this work, we propose a novel approach to unlearning in LLMs via soft prompting which is less resource-intensive than fine-tuning. To the best of our knowledge, this is the first work to investigate the use of soft prompting for unlearning in LLMs. Soft prompting simplifies the process of adapting LLMs to an arbitrary downstream task by optimizing learnable token embeddings that encode signals from a corresponding dataset (Lester et al., 2021; Li and Liang, 2021). The soft prompts are trained end-to-end and essentially act as instructions for a frozen pre-trained LLM during inference. We leverage this ability to modulate LLM outputs using prompts and formulate Soft Prompting for Unlearning (SPUL), a resource-efficient mechanism to achieve LLM unlearning in text classification. We optimize a set of soft prompt parameters that learn to encode underlying information in the data relevant for unlearning. When prepended to the input tokens of an LLM during inference, the soft prompts guide the LLM towards a generic response. We implement a multi-objective loss aligned with specific unlearning goals to facilitate the learning of soft prompts. SPUL unlearns undesirable outcomes without updating large-scale LLM parameters and can fully capitalize on the language understanding capability offered by the pre-trained LLMs. Consequently, our framework can utilize the same pre-trained LLM for different unlearning tasks and datasets during inference.
We evaluate SPUL for sentiment classification on benchmark NLP datasets when unlearning a subset from the corresponding training dataset and compare against various fine-tuning-based methods. We show that SPUL can effectively induce forgetting during inference while preserving the pre-trained utility with significant improvements over baselines. We conduct experiments to analyze the influence of SPUL hyperparameters including the contribution of loss components and the size of the soft prompts. We further validate SPUL on multiple pre-trained LLMs of different parameter sizes and different sizes of unlearning sets.
@@ 3.2 Problem Formulation
Given a training dataset $D^{tr}$ that was observed during pre-training of $h_{\theta}$, we assume a forget set, $D^{tr}_{f}\subset D^{tr}$, as the data intended for forgetting/removal from $h_{\theta}$. Simultaneously, we define a retain set $D^{tr}_{r}=D^{tr}\setminus D^{tr}_{f}$ comprising the remaining samples. Then, the goal of unlearning is to forget the token sequences in $D^{tr}_{f}$ while maintaining inference utility on $D^{tr}_{r}$. For our work, we focus on the task of text classification and interpret unlearning as the forgetting of the predictive output token sequences $y_{i}\in D^{tr}_{f}$. Essentially, we de-correlate text features and their corresponding labels for the relevant forget samples but preserve the classification performance on the retain samples. To this end, we aim to design a soft prompting framework to obtain optimized prompt tokens that can guide the base model toward the forget and retain objectives. With our framework, we aim to address the following research questions.
RQ1: How can soft prompting be utilized to effectively unlearn subsets of training data in the text classification domain?
@@ 3.3 Method
As soft prompts can be trained to encode signals from a dataset with the purpose of adapting a pre-trained LLM to a specific downstream task, we anticipate that the strategy can also be utilized to encode relevant information from an unlearning dataset containing forget and retain samples. Here, we propose the framework SPUL that leverages soft prompting to obtain effective prompt tokens $\boldsymbol{\phi}$ from an unlearning dataset $D^{tr}$ for text classification. Since one of the unlearning objectives in our framework is to promote feature and text de-correlation for forget samples, we design a loss attuned to enforcing misclassification for the respective text inputs. Specifically, we force the model to associate each input forget text sequence with a generic output token instead of its true label. We construct a generic label set $\bar{Y}$ that is disjoint from the task labels and contains tokens such as neutral or unknown and define a loss over the forget inputs,
@@
where $l(\cdot)$ again represents the cross-entropy loss. $L_{r}$ ensures that the model’s utility on the retain set does not degrade with the addition of prompt tokens. We further constrain the predictive distribution of the base model such that $h_{\theta}(\{\boldsymbol{\phi},\mathbf{x}_{j}\})$ reflects $h_{\theta}(\mathbf{x}_{j})$ for any $\mathbf{x}_{j}\in D^{tr}_{r}$. We quantify this difference using KL divergence as
@@ 4.1 Experimental Setup
We evaluate SPUL on two standard NLP datasets SST-2 (Socher et al., 2013) and Yelp polarity (Zhang et al., 2015) for the task of sentiment classification. The datasets contain reviews with each text sequence being labeled as a positive or negative sentiment. To build a realistic unlearning scenario where unlearning requests from each user would likely include multiple related training samples, we preprocess the datasets to construct the forget and retain sets such that the forget samples are semantically similar to each other (Yelp) or refer to common entities (SST-2).
For SST-2, we first perform Named Entity Recognition to identify named personalities, select a specific set of entities, and sample all related reviews to form the forget set $D^{tr}_{f}$. The remaining reviews are consequently assigned to the retain set $D^{tr}_{r}$. We perform a similar partitioning using the selected entities on the test set to obtain $D_{f}^{te}$ and $D_{r}^{te}$. After preprocessing, the constructed sets $D^{tr}_{f}/D^{tr}_{r}/D^{te}_{f}/D^{te}_{r}$ contains $1425/46331/610/19855$ samples. For the Yelp polarity dataset, we perform k-means clustering with cosine distance on the training data to divide the reviews into semantically similar groups. We randomly select a subset of the clusters and group them to form the $D^{tr}_{f}$ and the rest as $D^{tr}_{r}$. We utilize the same cluster centers to infer cluster identities for the test data and form the sets $D_{f}^{te}$ and $D_{r}^{te}$ accordingly. For Yelp, $D^{tr}_{f}/D^{tr}_{r}/D^{te}_{f}/D^{te}_{r}$ includes $5081/95012/885/18089$ samples.
@@
We use LLaMA-2-7B Touvron et al. (2023) as the base LLM to evaluate our SPUL framework. We further validate the unlearning effectiveness of our method with OPT-1.3B Zhang et al. (2022) and LLaMA-2-13B Touvron et al. (2023). To ensure familiarization with the unlearning dataset, we fine-tune the base LLMs on the full training dataset $D^{tr}=D^{tr}_{f}\cup D^{tr}_{r}$ for 10 (2) epochs on SST-2 (Yelp) with a learning rate set to 0.0001 and context length to 1024 using QLoRA (Dettmers et al., 2023). We treat this fine-tuned version of the LLM as the base model for unlearning. As for the configurations of SPUL, we fix the learning rate at 0.0001 across all LLMs, datasets and vary prompt token length $p$ among {10, 20, 30, 40, 50}. We also vary the regularization parameters $\alpha$ as {0.1, 0.5, 1.0} and $\beta$ as {0.0, 0.1, 0.5, 1.0} [^1] We note that advanced approaches, e.g., utility function, Pareto-based, and constraint-based methods, can be potentially adopted to determine values of $\alpha$ and $\beta$.. We train our unlearning framework for a total of 10 epochs. As for baseline model specifications, we conduct a parameter search for the best learning rates and report the most competitive results after 1 epoch of training. All experiments are conducted on NVIDIA A100 GPUs with 40GB RAM and we report the evaluation metrics over a single run due to the resource-intensive nature of the experiments.
@@ 4.2 Experimental Results
We include our main results with LLaMA-2-7B in Table 1. We report performance metrics for the original pre-trained LLM denoted as Vanilla and the fine-tuned base model denoted as QLoRA. We notice that the Vanilla results are considerably poorer for SST-2 compared to Yelp which validates our setup of fine-tuning the original LLM on the datasets for memorization. We attribute the difference in utility to the fact that the text sequences in Yelp are significantly longer and provide more contextual information. Nonetheless, after fine-tuning with QLoRA, the LLM’s performance increases to similar margins for both datasets.
From Table 1, we observe that SPUL significantly reduces accuracy and F1 on $D^{tr}_{f}$ compared to QLoRA demonstrating forgetting efficiency. At the same time, the difference in utility between SPUL and QLoRA for $D^{tr}_{r}$ is minimal showing that our method can promote unlearning while also preserving inference utility. Moreover, the metrics for $D^{te}_{f}$ and $D^{te}_{r}$ reflect those reported for $D^{tr}_{f}$ and $D^{tr}_{r}$ showing that the soft prompts effectively impose unlearning constraints on samples unseen during training. We observe similar performance trends for Yelp. Although the performance drop for $D_{f}^{tr}$ and $D_{r}^{te}$ are not equally as large as SST-2, the forget utility with the learned tokens is significantly lesser in comparison to the base model. We conjecture that the additional context provided by descriptive Yelp reviews restricts the forgetting capacity of the LLM. Also point out that utility loss in retain sets is smaller than forget sets.
Furthermore, SPUL outperforms baseline methods by a significant margin; compared to GA and RL, which optimize model parameters based only on the $D_{f}^{tr}$, SPUL consistently preserves inference utility on the retain sets with comparable or even lower metrics on the forget set. GA+KL and GA+GD optimize model parameters based on both $D_{f}^{tr}$ and $D_{r}^{tr}$. However, GA+KL performs poor on both datasets. GA+GD performs especially well on SST-2 but fails to enhance forget quality on Yelp which has more descriptive reviews. The proposed SPUL framework can however attain effective unlearning with the least loss of model utility. Among the compared methods, SPUL achieves significantly better overall trade-offs between the contrasting unlearning objectives of performance degradation and utility preservation.
@@
We conduct a series of experiments to investigate the influence of the hyperparameters $\alpha$ and $\beta$ on the unlearning performance of the proposed SPUL framework and report the results in Table 2 for the SST-2 dataset. The hyperparameters control the influence of the retain set on the learned soft prompts via losses $\mathcal{L}_{r}$ and $\mathcal{L}_{kl}$ respectively. We fix the number of prompt tokens $p$ at 30 for all results and vary $\alpha$ in {0.1, 0.5, 1.0} and $\beta$ among {0.0, 0.1, 0.5, 1.0}. From Table 2, we observe that at a fixed $\alpha$, unlearning efficacy is fairly unaffected by the change in the value of $\beta$. Model utility on the retain set, however, slightly increases as $\beta$ increases from 0.0 to 0.5 as $\mathcal{L}_{kl}$ gets more significance in the overall loss. We generally observe the best retain performance at $\beta$ = 0.5. The value of $\alpha$ influences performance on both forget and retain sets; higher $\alpha$ values benefit retain performance by prioritizing utility preservation whereas lower $\alpha$ values improve unlearning efficacy.
@@
We additionally evaluate the unlearning efficacy of our SPUL on different LLMs. In particular, we purposely choose OPT-1.3B with fewer parameters and LLaMA-2-13B with almost double the parameters compared to LLaMA-2-7B. In addition to the unlearning efficacy, this study also evaluates the scalability of our SPUL framework. We fix the hyperparameters $\alpha$ and $\beta$ at 1 and $p$ at 30 and report the results for SST-2 in Table 4. We first observe that the Vanilla inference with OPT-1.3B model performs noticeably poorer than LLaMA-2-7B whereas LLaMA-2-13B significantly improves over the initial metrics. This may be attributed to the pre-trained models’ complexity which affects their generalization ability. We similarly perform fine-tuning using QLoRA to ensure the unlearning dataset has been memorized by the respective LLM. Moreover, SPUL can effectively achieve the forget and retain unlearning objectives as made evident by the low forget accuracy and F1 compared to the retain metrics that closely resemble the base model’s performance. The results also indicate that the larger the LLM, the better it adapts to the unlearning task in our SPUL framework.

### v2

@@ Abstract
The widespread popularity of Large Language Models (LLMs), partly due to their unique ability to perform in-context learning, has also brought to light the importance of ethical and safety considerations when deploying these pre-trained models. In this work, we focus on investigating machine unlearning for LLMs motivated by data protection regulations. In contrast to the growing literature on fine-tuning methods to achieve unlearning, we focus on a comparatively lightweight alternative called soft prompting to realize the unlearning of a subset of training data. With losses designed to enforce forgetting as well as utility preservation, our framework Soft Prompting for Unlearning (SPUL) learns prompt tokens that can be appended to an arbitrary query to induce unlearning of specific examples at inference time without updating LLM parameters. We conduct a rigorous evaluation of the proposed method and our results indicate that SPUL can significantly improve the trade-off between utility and forgetting in the context of text classification and question answering with LLMs. We further validate our method using multiple LLMs to highlight the scalability of our framework and provide detailed insights into the choice of hyperparameters and the influence of the size of unlearning data. Our implementation is available at https://github.com/karuna-bhaila/llm_unlearning.
@@ 1 Introduction
In this work, we propose a novel approach to unlearning in LLMs via soft prompting which is less resource-intensive than fine-tuning. To the best of our knowledge, this is the first work to investigate the use of soft prompting for unlearning in LLMs. Soft prompting simplifies the process of adapting LLMs to an arbitrary downstream task by optimizing learnable token embeddings that encode signals from a corresponding dataset (Lester et al., 2021; Li and Liang, 2021). The soft prompts are trained end-to-end and essentially act as instructions for a frozen pre-trained LLM during inference. We leverage this ability to modulate LLM outputs using prompts and formulate Soft Prompting for Unlearning (SPUL), a resource-efficient mechanism to achieve LLM unlearning in text classification and multiple-choice question answering (MCQA). We optimize a set of soft prompt parameters that learn to encode underlying information in the data relevant for unlearning. When prepended to the input tokens of an LLM during inference, the soft prompts guide the LLM towards a generic response. We implement a multi-objective loss aligned with specific unlearning goals to facilitate the learning of soft prompts. SPUL unlearns undesirable outcomes without updating large-scale LLM parameters and can fully capitalize on the language understanding capability offered by the pre-trained LLMs. Consequently, our framework can utilize the same pre-trained LLM for different unlearning tasks and datasets during inference.
We evaluate SPUL for sentiment classification and MCQA tasks on benchmark NLP datasets when unlearning a subset from the corresponding training dataset and compare against various fine-tuning-based methods. We show that SPUL can effectively induce forgetting during inference while preserving the pre-trained utility with significant improvements over baselines. We conduct experiments to analyze the influence of SPUL hyperparameters including the contribution of loss components and the size of the soft prompts. We further validate SPUL on multiple pre-trained LLMs of different parameter sizes and different sizes of unlearning sets.
@@ 3.2 Problem Formulation
Given a training dataset $D^{tr}$ that was observed during pre-training of $h_{\theta}$, we assume a forget set, $D^{tr}_{f}\subset D^{tr}$, as the data intended for forgetting/removal from $h_{\theta}$. Simultaneously, we define a retain set $D^{tr}_{r}=D^{tr}\setminus D^{tr}_{f}$ comprising the remaining samples. Then, the goal of unlearning is to forget the token sequences in $D^{tr}_{f}$ while maintaining inference utility on $D^{tr}_{r}$. For our work, we focus on the task of text classification and question answering and interpret unlearning as the forgetting of the predictive output token sequences $y_{i}\in D^{tr}_{f}$. Essentially, we de-correlate text features and their corresponding labels for the relevant forget samples but preserve the predictive performance on the retain samples. To this end, we aim to design a soft prompting framework to obtain optimized prompt tokens that can guide the base model toward the forget and retain objectives. With our framework, we aim to address the following research questions.
RQ1: How can soft prompting be utilized to effectively unlearn subsets of training data in the text classification/QA domain?
@@ 3.3 Method
As soft prompts can be trained to encode signals from a dataset with the purpose of adapting a pre-trained LLM to a specific downstream task, we anticipate that the strategy can also be utilized to encode relevant information from an unlearning dataset containing forget and retain samples. Here, we propose the framework SPUL that leverages soft prompting to obtain effective prompt tokens $\boldsymbol{\phi}$ from an unlearning dataset $D^{tr}$. Since one of the unlearning objectives in our framework is to promote feature and text de-correlation for forget samples, we design a loss attuned to enforcing incorrect predictions for the respective text inputs. Specifically, we force the model to associate each input forget text sequence with a generic output token instead of its true label. We construct a generic label set $\bar{Y}$ that is disjoint from the task labels and contains tokens such as neutral, unknown, or None and define a loss over the forget inputs,
@@
where $l(\cdot)$ again represents the cross-entropy loss. $L_{r}$ ensures that the model’s utility on the retain set does not degrade with the addition of prompt tokens. In addition to maintaining performance on the retain set, the model after unlearning should closely resemble the model before unlearning. In our framework, we constrain the predictive distribution of the model such that $h_{\theta}(\{\boldsymbol{\phi},\mathbf{x}_{j}\})$ reflects $h_{\theta}(\mathbf{x}_{j})$ for any $\mathbf{x}_{j}\in D^{tr}_{r}$. We quantify this difference using KL divergence as
@@
Table 1: Dataset Statistics
Dataset $D^{tr}_{f}$ $D^{tr}_{r}$ $D^{te}_{f}$ $D^{te}_{r}$
SST-2 1425 46331 610 19855
Yelp polarity 5081 95012 885 18089
WMDP+SciQ 900 12679 373 1000
@@ 4.1 Experimental Setup
We evaluate SPUL on standard NLP datasets SST-2 (Socher et al., 2013) and Yelp polarity (Zhang et al., 2015) for the task of sentiment classification and datasets WMDP Li et al. (2024) and SciQ Welbl et al. (2017) for the task of multiple-choice question answering. SST-2 and Yelp contain reviews with each text sequence being labeled as a positive or negative sentiment. SciQ consists of science exam questions about Physics, Chemistry, and Biology, among others in a four-way multiple-choice format where each answer choice is associated with symbols such as “A”, “B”, etc. and WMDP contains questions about hazardous knowledge in biosecurity, cybersecurity, and chemical security in the same format. To build a realistic unlearning scenario where unlearning requests from each user would likely include multiple related training samples, we preprocess the classification datasets to construct the forget and retain sets such that the forget samples are semantically similar to each other (Yelp) or refer to common entities (SST-2). For MCQA, we construct forget sets to unlearn potentially harmful information while retaining general science knowledge.
For SST-2, we first perform Named Entity Recognition to identify named personalities, select a specific set of entities, and sample all related reviews to form the forget set $D^{tr}_{f}$. The remaining reviews are consequently assigned to the retain set $D^{tr}_{r}$. We perform a similar partitioning using the selected entities on the test set to obtain $D_{f}^{te}$ and $D_{r}^{te}$. For Yelp, we perform k-means clustering with cosine distance on the training data to divide the reviews into semantically similar groups. We randomly select a subset of the clusters and group them to form the $D^{tr}_{f}$ and the rest as $D^{tr}_{r}$. We utilize the same cluster centers to infer cluster identities for the test data and form the sets $D_{f}^{te}$ and $D_{r}^{te}$ accordingly. For the MCQA task, we construct the forget sets from WMDP containing questions about hazardous knowledge in biosecurity and the retain sets from SciQ with general science questions. We refer to this dataset as WMDP+SciQ. The sizes of the constructed forget and retain sets are reported in Table 1.
@@
We use LLaMA-2-7B Touvron et al. (2023) as the base LLM to evaluate our SPUL framework. We further validate the unlearning effectiveness of our method with OPT-1.3B Zhang et al. (2022) and LLaMA-2-13B Touvron et al. (2023). To ensure familiarization with the unlearning dataset, we fine-tune the base LLMs on the full training dataset $D^{tr}=D^{tr}_{f}\cup D^{tr}_{r}$ for 10 epochs on SST-2, 2 epochs on Yelp, and 5 epochs on WMDP+SciQ with a learning rate set to 0.0001 and context length to 1024 using QLoRA (Dettmers et al., 2023). We treat this fine-tuned version of the LLM as the base model for unlearning. As for the configurations of SPUL, we fix the learning rate at 0.0001 across all LLMs, and datasets and vary prompt token length $p$ among {10, 20, 30, 40, 50}. We also vary the regularization parameters $\alpha$ as {0.1, 0.5, 1.0} and $\beta$ as {0.0, 0.1, 0.5, 1.0} [^1] We note that advanced approaches, e.g., utility function, Pareto-based, and constraint-based methods, can be potentially adopted to determine values of $\alpha$ and $\beta$.. We train our unlearning framework for a total of 10 epochs. As for baseline model specifications, we follow earlier works Yao et al. (2024a) and conduct a parameter search for the best learning rates to report the most competitive results after 1 epoch of training. All experiments are conducted on NVIDIA A100 GPUs with 40GB RAM and we report the evaluation metrics over a single run due to the resource-intensive nature of the experiments.
@@ 4.2 Experimental Results
We include our main results with LLaMA-2-7B in Table 2. We report performance metrics for the original pre-trained LLM denoted as Vanilla and the fine-tuned base model denoted as QLoRA. We notice that the Vanilla results are considerably poorer for SST-2 compared to Yelp which validates our setup of fine-tuning the original LLM on the datasets for memorization. We attribute the difference in utility to the fact that the text sequences in Yelp are significantly longer and provide more contextual information. Nonetheless, after fine-tuning with QLoRA, the LLM’s performance increases to similar margins for both datasets. QLoRA fine-tuning similarly improves LLM’s predictions for the WMDP+SciQ dataset.
From Table 2, we observe that SPUL significantly reduces accuracy and F1 on $D^{tr}_{f}$ compared to QLoRA demonstrating forgetting efficiency. At the same time, the difference in utility between SPUL and QLoRA for $D^{tr}_{r}$ is minimal showing that our method can promote unlearning while also preserving inference utility. Moreover, the metrics for $D^{te}_{f}$ and $D^{te}_{r}$ reflect those reported for $D^{tr}_{f}$ and $D^{tr}_{r}$ showing that the soft prompts effectively impose unlearning constraints on samples unseen during training. We observe similar performance trends for Yelp and WMDP+SciQ. Although the performance drop for $D_{f}^{tr}$ and $D_{r}^{te}$ in Yelp are not equally as large as SST-2, the forget utility with the learned tokens is significantly lesser in comparison to the base model. We conjecture that the additional context provided by descriptive Yelp reviews restricts the forgetting capacity of the LLM. Also point out that utility loss in retain sets is smaller than forget sets. We also notice that SPUL performs exceedingly well on the WMDP+SciQ with the highest differences between the retain and forget metrics.
Furthermore, SPUL outperforms baseline methods by a significant margin; compared to GA and RL, which optimize model parameters based only on the $D_{f}^{tr}$, SPUL consistently preserves inference utility on the retain sets with comparable or even lower metrics on the forget set. For WMDP+SciQ, both baselines underperform in the forgetting task. GA+KL and GA+GD optimize model parameters based on both $D_{f}^{tr}$ and $D_{r}^{tr}$. However, GA+KL performs poorly on all three datasets. GA+GD performs especially well on SST-2 and WMDP+SciQ but fails to enhance forget quality on Yelp which has more descriptive reviews than SST-2. The proposed SPUL framework can however attain effective unlearning with the least loss of model utility. Among the compared methods, SPUL achieves significantly better overall trade-offs between the contrasting unlearning objectives of performance degradation and utility preservation for both tasks.
@@
Ablation Study
We conduct a series of experiments to investigate the influence of $\mathcal{L}_{r}$ and $\mathcal{L}_{kl}$ on the unlearning performance of the proposed SPUL framework and report the results in Table 3 for the SST-2 dataset. The hyperparameters $\alpha$ and $\beta$ control the influence of the retain set on the learned soft prompts via losses $\mathcal{L}_{r}$ and $\mathcal{L}_{kl}$ respectively. We fix the number of prompt tokens $p$ at 30 for all results and vary $\alpha$ in {0.1, 0.5, 1.0} and $\beta$ among {0.0, 0.1, 0.5, 1.0}. From Table 3, we observe that at a fixed $\alpha$, unlearning efficacy is fairly unaffected by the change in the value of $\beta$. Model utility on the retain set, however, slightly increases as $\beta$ increases from 0.0 to 0.5 as $\mathcal{L}_{kl}$ gets more significance in the overall loss. We generally observe the best retain performance at $\beta$ = 0.5. The value of $\alpha$ influences performance on both forget and retain sets; higher $\alpha$ values benefit retain performance by prioritizing utility preservation whereas lower $\alpha$ values improve unlearning efficacy.
@@
We additionally evaluate the unlearning efficacy of our SPUL on different LLMs. In particular, we purposely choose OPT-1.3B with fewer parameters and LLaMA-2-13B with almost double the parameters compared to LLaMA-2-7B. In addition to the unlearning efficacy, this study also evaluates the scalability of our SPUL framework. We fix the hyperparameters $\alpha$ and $\beta$ at 1 and $p$ at 30 and report the results for SST-2 in Table 5. We first observe that the Vanilla inference with OPT-1.3B model performs noticeably poorer than LLaMA-2-7B whereas LLaMA-2-13B significantly improves over the initial metrics. This may be attributed to the pre-trained models’ complexity which affects their generalization ability. We similarly perform fine-tuning using QLoRA to ensure the respective LLM has memorized the unlearning dataset. Moreover, SPUL can effectively achieve the forget and retain unlearning objectives as made evident by the low forget accuracy and F1 compared to the retain metrics that closely resemble the base model’s performance. The results also indicate that the larger the LLM, the better it adapts to the unlearning task in our SPUL framework. Additionally, our results with OPT-1.3B indicate that SPUL significantly outperforms all baseline methods for both retain and forget metrics. We note that we could not run experiments on baselines with LLaMA-2-13B due to limited GPU as all baseline methods require full fine-tuning of the LLM. This further highlights the advantage of SPUL over baselines in terms of parameter efficiency.